# Recent Advances in Large Langauge Model Benchmarks against Data Contamination: From Static to Dynamic Evaluation

### v1

@@ Abstract
Data contamination has received increasing attention in the era of large language models (LLMs) due to their reliance on vast Internet-derived training corpora. To mitigate the risk of potential data contamination, LLM benchmarking has undergone a transformation from static to dynamic benchmarking. In this work, we conduct an in-depth analysis of existing static to dynamic benchmarking methods aimed at reducing data contamination risks. We first examine methods that enhance static benchmarks and identify their inherent limitations. We then highlight a critical gap—the lack of standardized criteria for evaluating dynamic benchmarks. Based on this observation, we propose a series of optimal design principles for dynamic benchmarking and analyze the limitations of existing dynamic benchmarks. This survey provides a concise yet comprehensive overview of recent advancements in data contamination research, offering valuable insights and a clear guide for future research efforts. We maintain a GitHub repository to continuously collect both static and dynamic benchmarking methods for LLMs. The repository can be found at this link  [^1] GitHub Link: Static-to-Dynamic LLM Benchmarking.
Recent Advances in Large Langauge Model Benchmarks against Data Contamination: From Static to Dynamic Evaluation
Simin Chen1 Yiming Chen2 Zexin Li3 Yifan Jiang4 Zhongwei Wan5 Yixin He4 Dezhi Ran6 Tianle Gu7 Haizhou Li2,8 Tao Xie6 Baishakhi Ray1 1Columbia University 2National University of Singapore 3University of California, Riverside 4University of Southern California 5The Ohio State University 6Peking University 7Tsinghua University 8The Chinese University of Hong Kong, Shenzhen
@@ 1 Introduction
The field of natural language processing (NLP) has advanced rapidly in recent years, fueled by breakthroughs in Large Language Models (LLMs) such as GPT-4, Claude3, and DeepSeek Achiam et al. (2023); Liu et al. (2024); Wan et al. (2023). Trained on vast amounts of Internet-sourced data, these models have demonstrated remarkable capabilities across various applications, including code generation, text summarization, and mathematical reasoning Codeforces (2025); Hu et al. (2024).
@@
To develop and enhance LLMs, beyond advancements in model architectures and training algorithms, a crucial area of research focuses on effectively evaluating their intelligence. Traditionally, LLM evaluation has relied on static benchmarking, which involves using carefully curated human-crafted datasets and assessing model performance with appropriate metrics Wang (2018); Achiam et al. (2023); Gunasekar et al. (2023). For example, in 2021, OpenAI introduced the HumanEval Chen et al. (2021) dataset to benchmark the code generation capabilities of LLMs; HumanEval has since become a widely used standard for this purpose.
@@
However, because these static benchmarks are released on the Internet for transparent evaluation, and LLMs gather as much data as possible from the Internet for training, potential data contamination is unavoidable Magar and Schwartz (2022); Deng et al. (2024b); Li et al. (2024d); Sainz et al. (2024); Balloccu et al. (2024). Data contamination occurs when benchmark data is inadvertently included in the training phase of language models, leading to an inflated and misleading assessment of their performance. While this issue has been recognized for some time—stemming from the fundamental machine learning principle of separating training and test sets—it has become even more critical with the advent of LLMs. These models often scrape vast amounts of publicly available data from the Internet, significantly increasing the likelihood of contamination Achiam et al. (2023); Liu et al. (2024). Furthermore, due to privacy and commercial concerns, tracing the exact training data for these models is challenging, if not impossible, complicating efforts to detect and mitigate potential contamination.
Figure 1: The progress of benchmarking LLM
@@
To mitigate the risk of data contamination in LLM benchmarking, researchers have proposed various methods to enhance static approaches, such as data encryption Jacovi et al. (2023) and post-hoc contamination detection Shi et al. (2024). However, due to the inherent limitations of static methods, researchers have introduced new dynamic benchmarking schema (Fig. 1 illustrates the evolution from static to dynamic benchmarking). One of the dynamic benchmarking methods involves continuously updating benchmark datasets based on the timestamps of LLM training to minimize contamination White et al. (2024); Jain et al. (2024). Another method focuses on regenerating benchmark data to reconstruct original benchmarks Chen et al. (2024); Zhou et al. (2025); Mirzadeh et al. (2025), thereby reducing the likelihood of contamination.
Despite the numerous approaches proposed for transparent LLM benchmarking, no systematic survey has summarized the rapid progress of dynamic benchmarking methods. Existing literature reviews primarily focus on post-hoc detection of data contamination and do not address emerging dynamic benchmarking strategies Deng et al. (2024a); Ravaut et al. (2024). Moreover, no existing work discusses criteria for evaluating dynamic benchmarks themselves. Additionally, these reviews often lack a comprehensive discussion of the strengths and limitations of each method.
@@
To bridge this gap, we first conduct a systematic survey of benchmarking methods for LLMs designed to mitigate the risk of data contamination, covering both static and dynamic benchmarks. We summarize state-of-the-art methods and provide an in-depth discussion of their strengths and limitations. Furthermore, we are the first to propose a set of criteria for evaluating dynamic benchmarks. Our study reveals that existing dynamic benchmarks do not fully satisfy these proposed criteria, implying the imperfection of current design. We hope that our criteria will provide valuable insights for the future design and standardization of dynamic benchmarking methods.
Figure 2: Taxonomy of research on benchmarking LLMs
The paper is structured as illustrated in Fig. 2. We begin with an overview of data contamination (§2), establishing the necessary background. Next, we provide a comprehensive review of static benchmarking methods and examine how existing research has improved them to mitigate potential data contamination (§3). We then outline key criteria for an effective dynamic benchmark and introduce existing dynamic benchmarking approaches (§4). Finally, in §5, we reflect on insights from prior work, summarize current challenges, and explore future research directions.
@@
Table 1: Examples of Data Contamination in LLMs
@@ 2.1 Data Contamination
Data contamination occurs when LLM’s training data $\mathcal{D}_{\text{train}}$ contains information that improperly overlaps with evaluation benchmark data $\mathcal{D}_{\text{test}}$, compromising the validity of performance measurements. We summarize existing work and provide a formal definition of data contamination.
Exact contamination occurs when there is any exact duplicate in the benchmark dataset
@@
In other word, there exist a data point $d$ that both in $\mathcal{D}_{\text{train}}$ and $\mathcal{D}_{\text{test}}$. Common cases include verbatim test examples appearing in training corpora, code snippets from benchmark implementations, or documentation leaks.
Syntactic contamination occurs when a test data point could be found in the training dataset after a syntactic transformation, such that
@@
Examples of each contamination
We provide contamination examples in Table 1. In the case of syntactic contamination, the test data is derived from the training data by rephrasing it with the addition of a prefix string. There is ongoing debate about whether such syntactic transformations constitute true data contamination, as it is challenging to distinguish between an LLM’s ability to recall memorized information and its reasoning capability during inference. In this work, we consider such syntactic transformations as contamination, given that some NLP applications rely primarily on syntactic information for decision-making.
Significance of contamination
Understanding and mitigating potential data contamination in benchmarking LLMs is significant, especially given the rapid pace of LLM development. Without a robust approach to identifying and preventing contamination, evaluations may overestimate a model’s true capabilities by inadvertently testing it on data it has already seen. This undermines the validity of benchmarks, making it difficult to assess generalization, robustness, and real-world applicability. Contaminated benchmarks can lead to misleading conclusions about progress in LLM research, influencing model comparisons, deployment decisions, and policy-making. Addressing this issue is crucial for ensuring that benchmarks provide an accurate and reliable measure of an LLM’s true ability to handle novel and unseen data.
2.2 Contamination from LLM Training
Unlike traditional models with clear separations between training and evaluation data, LLMs are pre-trained on massive, diverse datasets—often scraped from the web (e.g., FineWeb Penedo et al. (2024))—which increases the risk of evaluation data overlap. In the post-training phase, models are further fine-tuned on large human-annotated Mukherjee et al. (2023); Kim et al. (2023) or synthetic datasets Ding et al. (2023); Teknium (2023); Wang et al. (2023) that may resemble evaluation tasks, further compounding contamination risks. Although retrieval-based detection methods Team et al. (2024); Achiam et al. (2023) exist, the sheer scale and complexity of training corpora make it difficult to entirely exclude evaluation data. Additionally, many LLMs keep their training data proprietary Dubey et al. (2024); Yang et al. (2024), complicating the accurate assessment of their true performance and highlighting the need for fair and reliable benchmarks. This opacity further exacerbates data contamination, as it impedes the community’s ability to verify and mitigate potential overlaps between training and evaluation data.
@@ 2.3 LLM Benchmarking
Task Type Benchmark
Language Static GLUE (Wang, 2018), SuperGLUE (Wang et al., 2019), CLUE (Xu et al., 2020)
Dynamic LiveBench (White et al., 2024), C2LEVA (Li et al., 2024c), ITD (Zhu et al., 2024c)
Dynamic LiveBench (White et al., 2024), LiveCodeBench (Jain et al., 2024), ComplexCodeEval (Feng et al., 2024)
Reasoning Static PIQA (Bisk et al., 2020), SIQA (Sap et al., 2019), HellaSwag (Zellers et al., 2019), WinoGrande (Sakaguchi et al., 2021), ARC (Clark et al., 2018), OpenBookQA (Mihaylov et al., 2018), CommonsenseQA (Talmor et al., 2018), C-SimpleQA (He et al., 2024)
Dynamic LiveBench (White et al., 2024), DyVal (Zhu et al., 2024a), C2LEVA (Li et al., 2024c), NPHardEval (Fan et al., 2024), S3Eval (Lei et al., 2024), DARG (Zhang et al., 2024b)
Knowledge Static NaturalQuestions (Kwiatkowski et al., 2019), TriviaQA (Joshi et al., 2017), CMMLU (Li et al., 2023a), MMLU (Hendrycks et al., 2020), BBH (Suzgun et al., 2022), AGI Eval (Zhong et al., 2023), MMLU-Redux (Gema et al., 2024), MMLU-Pro (Wang et al., 2024b), ControlBench (Darioush et al., 2024), FRAMES (Krishna et al., 2024), GPQA Diamond (Rein et al., 2023), AlpacaEval (Li et al., 2023c), ArenaHard (Li et al., 2024a)
Dynamic C2LEVA (Li et al., 2024c), ITD (Zhu et al., 2024c), Auto-Dataset (Ying et al., 2024), DyVal2 (Zhu et al., 2024b), SciEval (Sun et al., 2024)
Dynamic C2LEVA (Li et al., 2024c), FactBench (Bayat et al., 2024)
Dynamic LiveBench (White et al., 2024)
Dynamic LatestEval (Li et al., 2023d), Antileak-bench (Wu et al., 2024)
Table 2: Summary of benchmarking applications.
@@ 3 Static Benchmarking
In this section, we summarize a collection of static benchmarks that have been widely used to evaluate various aspects of model performance. These benchmarks cover a broad range of tasks including math, language, coding, reasoning, knowledge, safety, instruction following, and reading comprehension. They serve as standardized evaluation tools to measure model abilities in areas such as arithmetic problem-solving, natural language understanding, program synthesis, commonsense reasoning, factual knowledge retrieval, toxicity detection, and more. Table 2 provides an overview of these benchmarks along with the corresponding task categories and key references.
@@ 3.1 Problem Formulation
3.2 Static Benchmark Application
Math
@@ 3.2 Static Benchmark Application
Math benchmarks evaluate a model’s ability to solve multi-step math problems. Datasets such as GSM8K (Cobbe et al., 2021) and MATH (Hendrycks et al., 2021) require models to work through complex problems. Recent challenges like AIME 2024 (of America, 2024) and CNMO 2024 (Society, 2024) further test a model’s capacity to tackle diverse and intricate math tasks.
Knowledge
Knowledge benchmarks evaluate LLM internal knowledge. NaturalQuestions (Kwiatkowski et al., 2019) and TriviaQA (Joshi et al., 2017) focus on retrieving real-world information, while multi-domain tasks are covered by MMLU (Hendrycks et al., 2020), BBH (Suzgun et al., 2022), and AGI Eval (Zhong et al., 2023). Recent extensions like MMLU-Redux (Gema et al., 2024) and MMLU-Pro (Wang et al., 2024b) refine these assessments further. Additionally, ControlBench (Darioush et al., 2024), FRAMES (Krishna et al., 2024), and GPQA Diamond (Rein et al., 2023) target technical and long-context challenges, with open-domain evaluations provided by AlpacaEval (Li et al., 2023c) and ArenaHard (Li et al., 2024a).
Coding
@@
Coding benchmarks measure a model’s ability to generate and debug code. HumanEval (Chen et al., 2021) and MBPP (Austin et al., 2021) test code synthesis and debugging, whereas SWE-Bench (Jimenez et al., 2024; Yang et al., 2025) addresses more advanced challenges. Competitive platforms like Codeforces (Codeforces, 2025) and datasets such as Aider (Aider, 2025) further probe dynamic problem solving.
Instruction Following
@@
Instruction benchmarks evaluate a model’s ability to comprehend and execute detailed directives. Datasets like IFEval (Zhou et al., 2023) and InfoBench (Qin et al., 2024) simulate real-world scenarios requiring clear, step-by-step guidance, with C-Eval (Huang et al., 2024) focusing on Chinese instructions.
Reasoning
Understanding and applying everyday knowledge is a key aspect of language comprehension. Benchmarks such as PIQA (Bisk et al., 2020), SIQA (Sap et al., 2019), HellaSwag (Zellers et al., 2019), and WinoGrande (Sakaguchi et al., 2021) are designed to assess a model’s intuitive reasoning skills from multiple perspectives. In addition, academic challenge sets like ARC (Clark et al., 2018), OpenBookQA (Mihaylov et al., 2018), and CommonsenseQA (Talmor et al., 2018) push models further by requiring the integration of background knowledge with logical reasoning to arrive at plausible answers. C-SimpleQA (He et al., 2024) evaluates the factuality ability of language models to answer short questions in Chinese.
Safety
Safety benchmarks are essential for evaluating the robustness of LLM’s ability to generate non-toxic and ethically aligned content. Datasets such as RealToxicityPrompts (Gehman et al., 2020) and ToxiGen (Hartvigsen et al., 2022) assess resilience against producing harmful outputs. By providing a controlled environment to measure these aspects, safety benchmarks play a critical role in guiding the development of models that are not only powerful outputs but also responsible and trustworthy for real-world applications.
Language
Language benchmarks assess the LLMs’ proficiency in specific languages. GLUE (Wang, 2018) and SuperGLUE (Wang et al., 2019) cover tasks from sentiment analysis to language inference, while CLUE (Xu et al., 2020) targets Chinese language. Typo-fixing (Suzgun et al., 2022) is also widely used.
Reading Comprehension
Reading comprehension tasks test a model’s ability to extract and infer information from text. Benchmarks like SQuAD (Rajpurkar et al., 2018), QuAC (Choi et al., 2018), and BoolQ (Clark et al., 2019) challenge models to understand passages and draw logical conclusions.
3.3 Methods for Mitigation
Due to the nature of LLM training data collection and the public availability of these static benchmark datasets, there is a risk that LLMs may inadvertently encounter and use them, leading to data contamination. To address this issue, several methods have been proposed to enhance static benchmarking and mitigate the impact of data contamination.
3.3.1 Canary String
Canary strings are deliberately crafted, unique tokens embedded within a dataset to serve as markers for data contamination. When a model’s output unexpectedly includes these tokens, it strongly indicates that the model has memorized portions of its training data rather than learning to generalize. For instance, the BIG-Bench dataset incorporates these strings so that model developers can identify and filter out such instances Jacovi et al. (2023).
Limitations
The effectiveness of canary strings depends on model trainers being aware of and responsive to these markers. If a developer aims to leak benchmarking data to boost scores, this method will not work.
3.3.2 Encryption
Encryption methods secure evaluation data by making it inaccessible to unauthorized parties, preventing its accidental inclusion in training sets. Jacovi et al. (2023) propose encrypting test data with a public key and a “No Derivatives” license to block automated crawling and reuse. Yang et al. (2023) show that even advanced decontamination methods can be defeated by minor text variations, emphasizing the need for robust encryption. Similarly, TRUCE (Chandran et al., 2024) leverages confidential computing and secure multi-party computation to enable private benchmarking, ensuring that test data and model parameters remain confidential.
Limitation
While these methods effectively protect against data leakage, they depend on strong key management, they introduce extra computational overheads. These methods are vulnerable if encryption is compromised or private key is exposed.
3.3.3 Label Protection
Label protection involves keeping the true answers of a test set hidden from public access so that only an authorized evaluator can use them during model assessment. This approach is common in benchmarks such as GLUE Wang (2018), SuperGLUE Wang et al. (2019), and OpenAI’s HumanEval Chen et al. (2021), etc., where the test labels are withheld to prevent models from learning or memorizing them during training. The key advantage of this method is its ability to maintain evaluation integrity by preventing model exposure to answers, thereby mitigating data contamination risks.
Limitations
Label protection limits transparency and independent verification, and it forces researchers to rely on centralized evaluation systems for performance metrics, which can impede detailed error analysis and reproducibility.
3.3.4 Post-hoc Detection
Post-hoc detection mitigates data contamination by identifying overlaps between $D_{train}$ and $D_{test}$. This is typically done through n-gram matching at various levels, such as tokens Touvron et al. (2023) or words Radford et al. (2019); Brown et al. (2020); Chowdhery et al. (2023). However, exact matching often leads to false negatives, prompting the use of more robust techniques like embedding-based similarity Riddell et al. (2024); Lee et al. (2023); Gunasekar et al. (2023) and improved mapping metrics Li et al. (2024d); Xu et al. (2024).
Limitations
Post-hot detection methods face several limitations. Full access to the training dataset is often restricted due to legal and privacy constraints, making overlap detection challenging. Additionally, assumptions about model behavior, such as higher memorization or lower perplexity for contaminated instances, may not hold across different models and tasks.

### v2

# Benchmarking Large Language Models Under Data Contamination: A Survey from Static to Dynamic Evaluation

@@ Abstract
In the era of evaluating large language models (LLMs), data contamination has become an increasingly prominent concern. To address this data contamination risk, LLM benchmarking has evolved from a static to a dynamic paradigm. In this work, we conduct an in-depth analysis of existing static and dynamic benchmarks for evaluating LLMs. We first examine methods that enhance static benchmarks and identify their inherent limitations. We then highlight a critical gap—the lack of standardized criteria for evaluating dynamic benchmarks. Based on this observation, we propose a series of optimal design principles for dynamic benchmarking and analyze the limitations of existing dynamic benchmarks. This survey provides a concise yet comprehensive overview of recent advancements in data contamination research, offering valuable insights and a clear guide for future research efforts. We maintain a GitHub repository to continuously collect both static and dynamic benchmarks for LLMs.
Benchmarking Large Language Models Under Data Contamination: A Survey from Static to Dynamic Evaluation
Simin Chen1   Yiming Chen2thanks: Corresponding authors: yiming.chen@u.nus.edu, zli536@ucr.edu   Zexin Li311footnotemark: 1   Yifan Jiang4   Zhongwei Wan5   Yixin He4 Dezhi Ran6Tianle Gu7Haizhou Li2,8Tao Xie6Baishakhi Ray1 1Columbia University  2National University of Singapore  3University of California, Riverside 4University of Southern California  5The Ohio State University  6Peking University 7Tsinghua University  8The Chinese University of Hong Kong, Shenzhen Static-to-Dynamic-LLMEval GitHub Repository
@@ 1 Introduction
The field of natural language processing (NLP) has advanced rapidly in recent years, driven by breakthroughs in Large Language Models (LLMs) such as GPT-4, Claude3, and DeepSeek Achiam et al. (2023); Liu et al. (2024); Wan et al. (2023). Trained on vast amounts of Internet-sourced data, these models have demonstrated remarkable capabilities across various applications, including code generation, text summarization, computer use, and mathematical reasoning Codeforces (2025); Ran et al. (2024); Hu et al. (2024).
@@
To develop and improve LLMs, beyond advancements in model architectures and training algorithms, a crucial area of research focuses on effectively evaluating their intelligence. Traditionally, LLM evaluation has relied on static benchmarking, which involves using carefully curated human-crafted datasets and assessing model performance with appropriate metrics Wang et al. (2018); Achiam et al. (2023); Gunasekar et al. (2023); Ran et al. (2025).
@@
However, because these static benchmarks are released on the Internet for transparent evaluation, and LLMs gather as much data as possible from the Internet for training, potential data contamination is unavoidable Magar and Schwartz (2022); Deng et al. (2024b); Li et al. (2024d); Balloccu et al. (2024). Data contamination occurs when benchmark data is inadvertently included in the training phase of LLMs, leading to inflated and misleading performance assessments. Although this issue has long been recognized—rooted in the fundamental machine learning principle of separating training and test sets—it has become more critical with the rise of LLMs, which often scrape vast amounts of publicly available Internet data Achiam et al. (2023), increasing the risk of contamination. Furthermore, due to privacy and commercial concerns, tracing the exact training data of these models is challenging—if not impossible—complicating efforts to detect and mitigate potential contamination.
Figure 1: The progress of benchmarking LLMs.
@@
To mitigate the risk of data contamination in LLM benchmarking, researchers have proposed several enhancements to static evaluation methods, including data encryption Jacovi et al. (2023) and post-hoc contamination detection Shi et al. (2024). However, due to the inherent limitations of static methods—such as unverifiable data exposure—these enhancements have seen limited adoption. As a result, researchers have shifted toward new dynamic benchmarking paradigms, as illustrated in Fig. 1. Dynamic methods aim to reduce contamination risk either by continuously updating benchmark datasets based on LLM training timestamps White et al. (2024); Jain et al. (2024), or by regenerating test data to reconstruct and replace original benchmarks Chen et al. (2024a); Zhou et al. (2025); Mirzadeh et al. (2025).
Although many dynamic benchmarking methods have been proposed to promote fair and transparent evaluation of LLMs, most existing work primarily highlights the advantages of these dynamic benchmarks White et al. (2024). However, the question remains: what are the potential trade-offs of using dynamic benchmarks to evaluate LLMs? The limitations of dynamic benchmarking—such as the computational overhead of continuous updates, and the need for reliable timestamp metadata—are not yet fully explored.
Moreover, existing surveys on LLM data contamination have mainly focused on post-hoc detection methods Deng et al. (2024a); Ravaut et al. (2024); Xu et al. (2024a); Dong et al. (2024); Balloccu et al. (2024), offering little attention to the emerging landscape of dynamic benchmarking strategies. Considering the growing importance and adoption of dynamic benchmarking methods, it is essential to assess their effectiveness and limitations. Unfortunately, our empirical survey of existing dynamic benchmarking methods reveals that their evaluations are highly fragmented. To date, there is no systematic work that defines clear evaluation criteria for dynamic benchmarks themselves. Moreover, existing reviews often overlook a detailed comparison of the strengths and weaknesses of different dynamic methods, leaving a gap in understanding their practical trade-offs and applicability.
@@
To bridge this gap, we first conduct a systematic survey of benchmarking methods for LLMs designed to mitigate the risk of data contamination, covering both static and dynamic benchmarks. We summarize state-of-the-art methods and provide an in-depth discussion of their strengths and limitations. Furthermore, we are the first to summarize and abstract a set of criteria for evaluating dynamic benchmarks. Our study reveals that existing dynamic benchmarks do not fully satisfy these proposed criteria, implying the imperfection of current design. We hope that our criteria will provide valuable insights for the future design and standardization of dynamic benchmarking methods.
Figure 2: Taxonomy of research on benchmarking LLMs.
The paper is organized as shown in Fig. 2. We first review the background on data contamination (§2), and then survey static benchmarks and their enhancements for mitigating data contamination (§3). Next, we introduce key principles and existing methods for dynamic benchmarking (§4). Finally, we discuss open challenges and future directions (§5).
@@
Table 1: Examples of data contamination in LLMs.
@@ 2.1 Data Contamination
Data contamination arises when LLM training data $\mathcal{D}_{\text{train}}$ improperly overlaps with evaluation data $\mathcal{D}_{\text{test}}$, undermining performance validity. We review existing work and formalize the definition.
Exact Contamination. Exact contamination occurs when there is any exact duplicate in the benchmark dataset
@@
In other words, there exists a data point $d$ that is in both $\mathcal{D}_{\text{train}}$ and $\mathcal{D}_{\text{test}}$. Common cases include verbatim test examples appearing in training corpora, code snippets from benchmark implementations, or documentation leaks.
Syntactic Contamination. Syntactic contamination occurs when a test data point could be found in the training dataset after a syntactic transformation, such that
@@
Examples of Each Contamination. We provide contamination examples in Table 1. Syntactic contamination occurs when test data is rephrased from training data using a prefix. Whether such syntactic contamination constitutes true contamination is debated, as it is difficult to separate memorization from reasoning. In this work, we treat such transformations as contamination, since some NLP tasks rely heavily on syntax.
Significance of Contamination. Data contamination poses a serious threat to the integrity of LLM benchmarking, particularly as models grow in scale and are trained on vast publicly available corpora. Without proper safeguards, evaluations may inadvertently test models on data that they have seen during training, leading to inflated performance metrics and misleading claims about generalization and robustness. Recent studies underscore this concern: Schaeffer (2023) demonstrates that pretraining on test data can significantly distort evaluation outcomes; Balloccu et al. (2024) reveal how easily data contamination and evaluation malpractices can occur in closed-source LLMs; Xu et al. (2024b) propose methods to quantify such contamination; and Deng et al. (2024a) provide a comprehensive survey of existing risks and mitigation strategies. The issue gained public attention when Meta’s LLaMA 4 faced allegations of using a non-public version fine-tuned for benchmark gains Babic (2025), raising concerns about evaluation transparency—despite Meta’s denial of test set exposure. Such cases underscore the need for contamination-aware benchmarking to accurately assess LLM performance on truly unseen data. We also present a proof-of-concept evaluation in Appendix A to highlight the impact of data contamination.
2.2 Contamination Source
Data contamination can occur during the pre-training, post-training, or fine-tuning phases of LLM development. Unlike traditional models with clear separations between training and evaluation data, LLMs are pre-trained on massive, diverse datasets—often scraped from the web (e.g., FineWeb Penedo et al. (2024)), increasing the risk of evaluation data overlap. In the post-training phase, models are further fine-tuned on large human-annotated Mukherjee et al. (2023); Kim et al. (2023) or synthetic datasets Ding et al. (2023); Teknium (2023); Wang et al. (2023) that may resemble evaluation tasks, further compounding contamination risks. Although retrieval-based detection methods Team et al. (2024); Achiam et al. (2023) exist, the sheer scale and complexity of training corpora make it difficult to entirely exclude evaluation data. Additionally, many LLMs keep their training data proprietary Dubey et al. (2024); Yang et al. (2024), complicating the accurate assessment of their true performance and highlighting the need for fair and reliable benchmarks. This opacity further exacerbates data contamination, as it impedes the community’s ability to verify and mitigate potential overlaps between training and evaluation data.
@@ 3.1 Problem Formulation
3.2 Static Benchmark Applications
@@
Math. Math benchmarks evaluate a model’s ability to solve multi-step math problems. Datasets such as GSM8K (Cobbe et al., 2021) and MATH (Hendrycks et al., 2021) require models to work through complex problems. Recent challenges like AIME 2024 (of America, 2024) and CNMO 2024 (Society, 2024) further test a model’s capacity to tackle diverse and intricate math tasks.
@@
Coding. Coding benchmarks measure a model’s ability to generate and debug code. HumanEval (Chen et al., 2021) and MBPP (Austin et al., 2021) test code synthesis and debugging, whereas SWE-Bench (Jimenez et al., 2024; Yang et al., 2025) addresses more advanced challenges. Competitive platforms like Codeforces (Codeforces, 2025) and datasets like Aider (Aider, 2025) further probe dynamic problem solving. There are also benchmarks that evaluate LLMs’ ability to judge the correctness of code (Jiang et al., 2025).
@@
Instruction Following. Instruction benchmarks evaluate a model’s ability to comprehend and execute detailed directives. Datasets like IFEval (Zhou et al., 2023) and InfoBench (Qin et al., 2024) simulate real-world scenarios requiring clear, step-by-step guidance, with C-Eval (Huang et al., 2024) focusing on Chinese instructions.
Other Applications. We provide a detailed introduction to other applications in Appendix B, along with a further analysis on enhancing static benchmarks in Appendix C.